How can i convert this problem 19/20 in its other two form. I am doing percents thank you.
Answer by nerdybill(7384)

How can i convert this problem 19/20 in its other two forms:
.
Decimal form:
simply divide 19 by 20 to get 0.95
19/20 = 0.95
.
Percent form:
19/20 * 100 = 0.95 * 100 = 95%